1. A 150kg boulder is held 800m above the ground. 
  2. What is its potential energy? 
  3. What is its kinetic energy? 
  4. The boulder begins to fall. What is its potential energy when it is 200m above the ground? Where did the “lost” potential energy go? 
  5. What is the kinetic energy of the boulder when it has fallen 200m? 
1
Expert's answer
2021-08-02T11:10:20-0400

(a) The potential energy at 800 m

"E_{p1}=mgh\\
=\rm 150\: kg*9.8\: m/s^2*800\: m=1.176*10^6\: J"

(b) The kinetic energy at 800 m

"E_{k1}=\frac{mv^2}{2}=0\:\rm J"

(c) The potential energy at 200 m

"E_{p2}=mgh\\
=\rm 150\: kg*9.8\: m/s^2*200\: m=2.94*10^5\: J"

(d) The kinetic energy at 200 m

"E_{k2}=E_{p1}-E_{p2}-E_{k1}\\
=\rm 1.176*10^6\: J-2.94*10^5\: J-0\:\rm J=8.82*10^5\: J"